# Exploiting Kubernetes’ Image Pull Implementation to Deny Node Availability

### v1

@@ V-B Denial of Resources and Scheduling Delay
We focus on assessing the denial of resources capabilities of the attack. In particular, we investigate to what extent the Pod scheduling can be delayed and how the CPU and I/O are affected.
@@ V-B3 Results
Variable GB set. We first run the experiments of the Variable GB image set. Before each experiment, we reshuffle the images. Then, for each image, we deploy it, wait two seconds, and then delete it with the ---force flag. Additionally, to verify the attack’s behavior with the parallelism provided by the $mp$ flag, we run each experiment with the values described before. We repeat this process thirty times for each setting; the final result is an average of these runs. To better visualize the results, we chose a random trial from the thirty and plotted the CPU usage, broken down by type, and the disk write and network read. These plots are available in 1(a) and 1(b).
With $mp$ set to 1, the results are staggering. Even with a small attack, no parallelism and just seven images, the attack generates a significant impact on the cluster, which clearly can be categorized as a DoS. Every new download on the node is blocked for a total of 1352 seconds, almost 23 minutes; the CPU usage averages 65% with spikes of almost 90%, and the disk is kept constantly busy, with I/O write spikes of 150 MB/s. From 1(a) (row 1), we can see how these CPU peaks are noticeable and correlate with the download of the images (1(b)), and thus with a high bandwidth and I/O usage. On the other hand, when images are decompressed, the CPU usage drops significantly but remains steadier, as does the disk usage.
As we increase $mp$, the total time of the attack decreases, and the resource usage increases accordingly. For example, setting $mp$ to $2$ reduces the attack time to 909 seconds, while with $mp$ $=4$, it further reduces to 887 seconds. On the other hand, the CPU usage increases, while the difference in the peaks becomes less and less apparent as the various parallel downloads compete for CPU and I/O time with the ongoing unpacking tasks. With four parallel images, the CPU usage reaches almost 100% for most of the attack time.
By calculating the Scheduling Delay for these three tests, we see how it also drops from 46.82 ($mp$ $=1$) to 31.48 ($mp$ $=2$) and 30.72 ($mp$ $=4$) s/GB. It is important to observe the difference between the former two settings, where the SD drops 33%, and the latter, where the SD drops again but by a smaller margin (only 2.4%). While introducing parallelism strengthens the attack and puts an increased load on resources, at a certain point, the returns start diminishing as contention is introduced, and the cluster’s overall performance starts degrading. Thus, we believe that increasing $mp$ without implementing a solution to this threat could be even more detrimental.
@@
From the Scheduling Delay results, the immediate impression is that using smaller images is indeed more disruptive. With $mp$ $=1$, the Variable GB set hangs the cluster for 47 seconds per GB. On the other hand, the Variable MB manages to block it for 57 seconds. Thus, we can block the downloads for 18% more time with small images. With $mp$ $=2$ and $mp$ $=4$, this becomes more evident, arriving at 23% more with small images. We attribute this increase to the overall overhead that containerd must withstand when downloading images. Indeed, each new image implies fetching its manifest, obtaining the list of layers, and manually downloading each of them: all operations repeated many times start piling up.
@@
Finally, we repeat our experiments in the GKE deployment as outlined in subsection V-A, obtaining consistent results. As GKE has yet to introduce support for the MaxParallelImagePulls parameter, we only test the scenario with sequential downloads. The Scheduling Delay of the attack decreases due to the increased bandwidth available in the GKE cluster. Additionally, a slightly weaker CPU contributes to an overall increase in CPU usage during the attack. However, such slight differences do not change the validity of our previous observations.
@@ V-C Service Interference
Another goal of the attack is to affect the Quality of Service (QoS) of other running services in the node. We perform two experiments with the same attack, but this time, we verify how this affects CPU and I/O usage of other Pods.
To do so, we employ the Phoronix Test Suite[17], an automated benchmark suite that allows users to perform a wide range of performance assessments on their system. From the wide variety of benchmarks available, we choose the build-linux-kernel [^13] Specifications available at https://openbenchmarking.org/test/pts/build-linux-kernel-1.15.0. The “defconfig” option was used. and stress-ng [^14] Specifications available at https://openbenchmarking.org/test/pts/stress-ng. The “cpustress” option was used. benchmarks, two of the most popular of the suite. The first assesses the time needed to compile the Linux kernel (version 6.1) in the machine, putting an I/O and CPU workload on the system. The latter evaluates the amount of operations per second the CPU can perform. To ensure the statistical relevance of its results, Phoronix automatically performs additional trials if it encounters performance issues or suspiciously high variability between different runs.
First, we run the build-linux-kernel benchmark with no attack running on the cluster. In this case, it completes in 853.25 seconds, averaged over three runs. Then, we apply the attack as described in subsection V-B with the Variable GB set and $mp$ set to 1 (as parallelism is not a concern in this setting). We choose this particular scenario as it is the longest and, thus, the most likely to last the entirety of the benchmark. We therefore re-run it, applying the attack in the background and cleaning up after every trial. With this setup, the build time increases to 1630.04, a 91% increase over the non-disrupted one.
The other benchmark also reports significant results, with the stress-ng measuring 2185.83 bogo-ops/s [^15] Bogus operations per seconds, a measure used by stress-ng to evaluate processor performance with no attack running over three trials. Repeating it with the attack, Phoronix takes 15 trials to obtain 1463.18 bogo-ops/s, a drop of 33% with regards to the non-attacked scenario. Additionally, the standard error is 12.52 seconds, almost ten times higher than the 1.29 seconds of the non-attacked benchmark. This 33% drop is less severe than the one from build-linux-kernel, mainly due to the fact that stress-ng’s CPU-bound behavior clashes more with the attack, introducing additional contention.
To further confirm our findings, we replicate our experiment in the GKE public cloud scenario, in identical conditions. With this setup, the resource impact is still significant. The build-linux-kernel benchmark over 5 trials increases the build time by 80%, while stress-ng reports an average of 45% less bogo ops/s over 15 trials. Thus, the impact on the benchmarks run on a public cloud is evident and comparable to the one on private cloud. Since on this setup the CPU is slightly weaker and the bandwidth is greater, the images are downloaded faster, but the CPU is used more in the process. Thus, CPU-bound benchmarks such as stress-ng suffer a greater drop than on the original environment. On the other hand, the disk on the GKE VM is faster, resulting in a lesser impact on the build-linux-kernel benchmark.
@@ V-D Cache Manipulation
Finally, we verify the attack’s capabilities in interfering with another kind of system resource: storage. In particular, we analyze how the attack can manipulate the local image cache.
As previously described in subsection IV-C, K8s is equipped with a local image cache which is periodically expunged by a GC should the used disk space surpass the ImageGCHighThresholdPercent (85%) threshold. Thus, we aim to investigate if the attack can have the side effect of inducing cache evictions. Such scenario would mean that the attacker can have a way of indirectly controlling the cache. In addition, it would further disrupt the node’s operations as the I/O and CPU overhead of the cache eviction is added to the underlying attack.
We therefore run again the Variable GB experiment with $mp$ being set to 1, as we are not interested in seeing the effects of parallel downloads in this case. With a simple script, we monitor the images present in the cache and the disk space. However, this time, we run the attack sequentially rather than randomly to eliminate variability. In other words, first we deploy the smallest image (2 GB), delete it with ---force, deploy the second smallest, and so on. We repeat the experiment several times, obtaining varying results. An example result is shown in Figure 3.
@@
V-E Attack Feasibility
The results presented in the subsubsection V-B3 and subsection V-D are strongly related to the resources available in the node. That opens an important discussion point about the attack feasibility generalization. We understand that generally, K8s worker nodes will have a limited size, mainly because of the applications’ nature running on it — as they are stateless — which usually do not maintain any status. In addition, bigger nodes will commonly be linked with large applications. That incentivizes a more spread use of resources between distinct nodes (i.e., horizontal node autoscaling) instead of a resource increase in the nodes already deployed.
These scenarios are ideal for small resource-based DoS attacks in specific nodes, where a continuous flow of CPU and I/O resources can deny a given application or force the eviction to another node. The same could be pernicious in Edge nodes, where the constrained resources with no redundancy are highly vulnerable to this kind of attack. Finally, since the spread usage of the resources and expected applications’ burst usage, it is almost impossible, with the current tools, to mitigate the threat only with resource monitoring.
@@ VI-B MAGI System
As a stopgap and proof-of-concept solution, we propose the MAGI System (Mitigation AGainst Images) [18], whose architecture is shown in Figure 4. It employs eBPF, Python, and Bash and monitors the cluster status and image downloading requests, taking prompt action whether an attack matching our pattern is detected.
@@
This is done first by querying the bbolt database181818bbolt is a fast on-disk database used by containerd. and obtaining the list of layers from the name of the image. Alternatively, the component can also query the registry directly, obtaining the list of layers from the manifest file. From each layer in the list, the port number of the connection downloading that particular layer is obtained. Finally, each of the offending sockets is terminated using ss -K. Since the Kubernetes API has already deleted the Pod and freed up its resources, it will not attempt again to download the image, thwarting the malicious actor’s attack. Clearly, if the attacker were to try again to download said image (or any other one), the system would get triggered again, thus being able to continously take down attack attempts over time.
@@
VI-C MAGI Evaluation
To evaluate the effectiveness of our mitigation approach, we devised a series of three tests that evaluate the functionality of the system in different scenarios: without the attack, with the attack, and with the attack and the mitigation running.
@@ VI-C MAGI Evaluation
In the first scenario, the total time spent by the test is around eleven minutes. Of these, eight of them are spent downloading the cocalc image, which is extremely large; once the download is completed, the other images follow suit quickly. On the other hand, in the second scenario we deploy the attack after the first twenty seconds, thus queuing it after the cocalc download. Once cocalc finishes downloading, all the following requests are brought to a grinding halt as the attack starts blocking the cluster. This increases the length of the test to around twenty-nine minutes. It is important to note that the attack does not affect the net download time of the images; rather, it makes them stall and wait until their turn arrives. Finally, in the third scenario, the MAGI System successfully intercepts and blocks six out of seven images, i.e., all images but the smallest one. As the downloads get interrupted almost instantly, the total length of the test drops back to eleven minutes, the same as the first scenario.
@@
These tests highlighted the capabilities of our solution. First, we observed how MAGI has almost no footprint. The CPU usage in nodes spikes at 0.8% during detection and almost 0% when sitting idle. The memory footprint was more evident, ranging from 380 MB idle to almost 420 MB during the attack, but remaining mostly stable nevertheless. We attribute this usage to Python’s memory management and the need of storing status for several different metrics at once. On the other hand, both CPU and memory usage were insignificant on the master node, with the only discernible overhead being K8s’ auditing feature.
However, the fact that MAGI failed to block the smallest image of the attack further strengthens the need for a more systematic and structured approach to the problem. While the eBPF code always identifies all offending downloads, the process of tracing back to the original layer is cumbersome. With enough bandwidth, small images get downloaded so quickly that it is almost impossible to even attempt blocking their download. In our opinion, the sheer amount of information required for gracefully terminating a download without disrupting any service, the inherent unpredictability of the system scheduler and the relative opacity of containerd’s download management render any further attempts in increasing MAGI’s speed better spent in improving the main problem instead.
@@ VII Related Work
Denial of Service based on Resources Usage. Resource-based DoS in Cloud Computing was previously covered in previous works. In [23], the authors present a novel class of attacks called Resource-Freeing Attacks (RFAs) and create a methodology to prevent the DoS by increasing the amount of resources available to the victim VM. Fang et al. also discussed resource and co-location attacks in [24] and [25], where they proposed respectively one new metric, called Heteroscore and a threat method, REPTTACK and present a mitigation technique to this type of attack. Finally, in [26], Zhan et al. show a framework that aims to detect CPU-exhaustion DoS attacks in containers called Coda. However, it does not consider the node resources, just the container ones.
@@
Kubernetes and Container Security. Ahmed [27] shows that the container deployment process can generate significant resource usage, mainly memory and CPU. The author also tested several alterations to the container deployment’s decompress process, focusing on the parallelism of several simultaneous layers. Such approach decreases the total time to deploy a given image but inherently increases the resources used in this process. Kubernetes is also focused in two papers on vulnerabilities in the configuration of distinct services in the orchestrator. He and Guo [3] show a threat in the resource configuration, where using highly privileged configurations allows eBPF applications to escape containers and access sensible information. Xiao [4], instead, found several threats based on the communication between microVMs solutions — such as Firecracker and Kata Containers — and containerd to escape attacks and resource-based DoS.
@@
Another important area that has received recent attention from the container orchestration community is anomaly detection in cluster resource usage. Almaraz-Rivera [5]’s study highlights the importance of implementing monitoring and alerting systems to track containers’ performance and resource utilization. He and Guo also implement a demo solution using an ML auto-encoder algorithm. Lastly, Yolchuyev [28] proposes an eXtreme Gradient Boosting (XGBoost) based model for anomaly detection on the Kubernetes orchestration platform. Either case, to the best of our knowledge no prior work has been conducted on exploiting the asynchronous design of the K8s API communication.

### v2

@@ V-B Denial of Resources and Scheduling Delay
We first focused on assessing the denial of resources capabilities of the attack. In particular, we investigated to what extent the Pod scheduling can be delayed and how the CPU and I/O are affected. These tests were performed on the local testbed.
@@ V-B3 Results
Variable GB set. We first run the experiments of the Variable GB image set on the local testbed. Before each experiment, we reshuffle the images. Then, for each image, we deploy it, wait two seconds, and then delete it with the ---force flag. Additionally, to verify the attack’s behavior with the parallelism provided by the $mp$ flag, we run each experiment with the values described before. We repeat this process thirty times for each setting; the final result is an average of these runs. To better visualize the results, we chose a random trial from the thirty and plotted the CPU usage, broken down by type, and the disk write and network read. These plots are available in 2(a) and 2(b).
With $mp$ set to 1, the results are staggering. Even with a small attack, no parallelism and just seven images, the impact on both worker nodes is substantial, and clearly could be categorized as a DoS. Every new download on the node is blocked for a total of 1352 seconds, almost 23 minutes; the CPU usage averages 65% with spikes of almost 90%, and the disk is kept constantly busy, with I/O write spikes of 150 MB/s. From 2(a) (row 1), we can see how these CPU peaks are noticeable and correlate with the download of the images (2(b)), and thus with a high bandwidth and I/O usage. On the other hand, when images are decompressed, the CPU usage drops significantly but remains steadier, as does the disk usage.
As we increase $mp$, the total time of the attack decreases, and the resource usage increases accordingly. For example, setting $mp$ to $2$ reduces the attack time to 909 seconds, while with mpmp =4=4, it further reduces to 887 seconds. On the other hand, the CPU usage increases, while the difference in the peaks becomes less and less apparent as the various parallel downloads compete for CPU and I/O time with the ongoing unpacking tasks. With four parallel images, the CPU usage reaches almost 100% for most of the attack time.
By calculating the Scheduling Delay for these three tests, we see how it also drops from 46.82 ($mp$ =1=1) to 31.48 ($mp$ =2=2) and 30.72 ($mp$ =4=4) s/GB. It is important to observe the difference between the former two settings, where the SD drops 33%, and the latter, where the SD drops again but by a smaller margin (only 2.4%). While introducing parallelism strengthens the attack and puts an increased load on resources, at a certain point, the returns start diminishing as contention is introduced, and the cluster’s overall performance starts degrading. Thus, we believe that increasing $mp$ without implementing a solution to this threat could be even more detrimental.
@@
From the Scheduling Delay results, the immediate impression is that using smaller images is indeed more disruptive. With mpmp =1=1, the Variable GB set hangs the worker nodes for 47 seconds per GB. On the other hand, the Variable MB manages to block it for 57 seconds. Thus, we can block the downloads for 18% more time with small images. With mpmp =2=2 and mpmp =4=4, this becomes more evident, arriving at 23% more with small images. We attribute this increase to the overall overhead that containerd must withstand when downloading images. Indeed, each new image implies fetching its manifest, obtaining the list of layers, and manually downloading each of them: all operations repeated many times start piling up.
@@ V-C Service Interference
Another goal of the attack is to affect the Quality of Service (QoS) of other running services in the node. We performed two experiments on the local testbed with the same attack, but this time, we verified how this affects CPU and I/O usage of other Pods.
To do so, we employ the Phoronix Test Suite[18], an automated benchmark suite that allows users to perform a wide range of performance assessments on their system. From the wide variety of benchmarks available, we choose the build-linux-kernel [^10] Specifications available at https://openbenchmarking.org/test/pts/build-linux-kernel-1.15.0. The “defconfig” option was used. and stress-ng [^11] Specifications available at https://openbenchmarking.org/test/pts/stress-ng. The “cpustress” option was used. benchmarks, two of the most popular of the suite. The first assesses the time needed to compile the Linux kernel (version 6.1) on the machine, putting an I/O and CPU workload on the system. The latter evaluates the amount of operations per second the CPU can perform. To ensure the statistical relevance of its results, Phoronix automatically performs additional trials if it encounters performance issues or suspiciously high variability between different runs.
First, we run the build-linux-kernel benchmark with no attack running on the cluster. In this case, it completes in 853.25 seconds, averaged over three runs. Then, we apply the attack as described in subsection V-B with the Variable GB set and $mp$ set to 1 (as parallelism is not a concern in this setting). We chose this particular scenario as it is the longest and, thus, the most likely to last the entirety of the benchmark. We therefore re-run it, applying the attack in the background and cleaning up after every trial. With this setup, the build time increases to 1630.04, a 91% increase over the non-disrupted one.
The other benchmark also reports significant results, with the stress-ng measuring 2185.83 bogo-ops/s [^12] Bogus operations per second, a measure used by stress-ng to evaluate processor performance with no attack running over three trials. Repeating it with the attack, Phoronix takes 15 trials to obtain 1463.18 bogo-ops/s, a drop of 33% with regards to the non-attacked scenario. Additionally, the standard error is 12.52 seconds, almost ten times higher than the 1.29 seconds of the non-attacked benchmark. This 33% drop is less severe than the one from build-linux-kernel, mainly due to the fact that stress-ng’s CPU-bound behavior clashes more with the attack, introducing additional contention.
@@ V-D Cache Manipulation
We then verified the attack’s capabilities in interfering with another kind of system resource: storage. In particular, we analyze how the attack can manipulate the local image cache on the local testbed.
As previously described in subsection IV-D, K8s is equipped with a local image cache which is periodically expunged by a GC should the used disk space surpass the ImageGCHighThresholdPercent (85%) threshold. Thus, we aim to investigate whether the attack can induce cache evictions as a side effect. Such a scenario would mean that the attacker can have a way of indirectly controlling the cache. In addition, it would further disrupt the node’s operations as the I/O and CPU overhead of the cache eviction is added to the underlying attack.
We therefore rerun the Variable GB experiment with $mp$ set to 1, as we are not interested in seeing the effects of parallel downloads in this case. With a simple script, we monitor the images present in the cache and the disk space. However, this time, we run the attack sequentially rather than randomly to eliminate variability. In other words, first we deploy the smallest image (2 GB), delete it with ---force, deploy the second smallest, and so on. We repeat the experiment several times, obtaining varying results. An example result is shown in Figure 3.
@@
V-E Applicability to Public Cloud
To further confirm our findings, we conducted additional experiments in the GKE public cloud scenario, aiming to determine if the attack was feasible in larger and more realistic scenarios. First, we replicated all the experiments described so far in this section. The results were consistent with those obtained in the local testbed, but due to space constraints, they can be found in the repository. Second, to investigate whether increasing the number of nodes affects the attack’s capabilities, we crafted a new scenario using a set of real, Machine Learning (ML)-oriented images and a slightly different attack procedure. Additionally, information about the setup can be found in the repository.
In this scenario, instead of creating and deleting Pods, we first create a Deployment. The Deployment is configured to have eight replicas, one for each node in the cluster, using affinity features to achieve so. Since the images are large, we wait several seconds after the first creation, and then we patch the Deployment, changing the image. As a result, the underlying ReplicaSet is substituted, triggering a process in which old containers are deleted, and new ones are spawned with the new image, thus requiring a new download from the container registry. This process realizes the same effect as the previous attack with the ---force flag, but without actually using it. We also validate whether Falco can identify this new approach. Yet, with the default set of rules, the only INFO alert printed is the one related to the first Deployment creation, since ReplicaSets are not observed by Falco.
This attack proves, in a production scenario, that all the characteristics explored in the local testbed (the scheduling delay, resource usage, and cache manipulation) are still valid. With just a single alert triggered, an attacker could perpetually edit the Deployment, keeping the node(s) under constant pressure. The CPU usage over all nodes is still high, averaging 91% over the attack duration. Each node reacted differently to the various download requests: unfortunately, some images were never downloaded by some nodes. Yet, all of them were overall delayed for a 18 minutes on average. The GC could have also been triggered if we had increased the number of images. Thus, we can conclude that the attack is indeed feasible in a public cloud environment. Attackers wishing to make it stealthier could reduce the amount of replicas or use specific nodeSelectors, achieving a similar effect. The only required permission is the ability to create and patch Deployments, and the number of nodes that can be affected is limited by the amount of replicas one can create.
@@ VI-B MAGI System
As a stopgap and proof-of-concept solution, we propose the MAGI System (Mitigation AGainst Images) [19], whose architecture is shown in Figure 4. MAGI is fully deterministic and is written in eBPF, Python, and Bash. It monitors the cluster status and image downloading requests, and uses kernel-level tools to terminate offending image downloads whether an attack matching our pattern is detected.
@@
MAGI can be deployed in any K8s cluster, regardless of the container runtime being used. It requires minimal effort to be set up, with the only requirement being the enabling of K8s’ auditing feature and root access to the nodes. Once deployed, it runs in the background and does not require any user interaction.
@@
This is done first by querying the bbolt database151515bbolt is a fast on-disk database used by containerd. and obtaining the list of layers from the name of the image. Alternatively, the component can also query the registry directly, obtaining the list of layers from the manifest file. From each layer in the list, the port number of the connection downloading that particular layer is obtained. Finally, each of the offending sockets is terminated using ss -K. Since the Kubernetes API has already deleted the Pod and freed up its resources, it will not attempt again to download the image, thwarting the malicious actor’s attack. Clearly, if the attacker were to try again to download said image (or any other one), the system would get triggered again, thus being able to continously take down attack attempts over time.
@@
VI-C Evaluation
To evaluate the effectiveness of our mitigation approach, we devised a series of tests that evaluate the functionality of the system in different scenarios: without the attack, with the attack, and with the attack and the mitigation running.
@@ VI-C Evaluation
In the first scenario, the total time spent by the test is around eleven minutes. Of these, eight of them are spent downloading the cocalc image, which is extremely large; once the download is completed, the other images follow suit quickly. On the other hand, in the second scenario we deploy the attack after the first twenty seconds, thus queuing it after the cocalc download. Once cocalc finishes downloading, all the following requests are brought to a grinding halt as the attack starts saturating the CPU and bandwidth of the worker nodes. This increases the length of the test to around twenty-nine minutes. It is important to note that the attack does not affect the net download time of the images; rather, it makes them stall and wait until their turn arrives. Finally, in the third scenario, the MAGI System successfully intercepts and blocks six out of seven images, i.e., all images but the smallest one. As the downloads get interrupted almost instantly, the total length of the test drops back to eleven minutes, the same as the first scenario.
@@
These tests highlight the capabilities and low impact of our solution. First, we observed how MAGI has almost no footprint. The CPU usage in nodes spikes at 0.8% during detection and almost 0% when sitting idle. The memory footprint was more evident, ranging from 380 MB idle to almost 420 MB during the attack, but remaining mostly stable nevertheless. We attribute this usage to Python’s memory management and the need of storing status for several different metrics at once. On the other hand, both CPU and memory usage were insignificant on the master node, with the only discernible overhead being K8s’ auditing feature.
Wishing to further investigate why MAGI failed to block the smallest image of the attack, we performed a series of tests with a set of smaller images, starting from 83 MB (42.5 MB uncompressed) and adding a new 5 MB layer at each step. With MAGI running in the background, we deployed and deleted a Pod with the image, waited for the download to complete, purged the node’s image cache, and then repeated the process with the next image. We observed that the cutoff in which MAGI was able to block the download was around 250 MB, which on a 10 Gbit/s network takes around 2 seconds to download. MAGI’s eBPF code identified all image downloads and the corresponding layers, but with such a large bandwidth, downloads complete so quickly that is almost impossible to terminate the socket before the download finishes. With smaller bandwidths, the download time increases, shifting the size threshold accordingly.
VI-D Discussion and limitations
While MAGI demonstrates a significant improvement over the current state of affairs, it is does not come without drawbacks. First and foremost, the attack surface of the cluster increases significantly when MAGI is deployed. This is due to several factors.
First, the K8s auditing feature – which is not enabled by default – exposes sensitive information about the cluster’s state and operations. Second, MAGI needs to monitor the gRPC calls between the API server and the container runtime. Doing so requires restarting the Kubelet, which can disrupt the cluster’s operations. Finally, the MAGI client requires root access to the nodes in order to intercept the system calls and kill the offending sockets. All of these extra requirements increase the attack surface.
In addition to the security risks, MAGI also has some limitations in terms of performance and usability. As shown above, MAGI struggles in blocking the download of very small images. Indeed, the time required to perform heavy tasks such as the domain resolution, manifest retrieval, and socket termination can exceed the time required to download the image itself, if it is small. MAGI is also inherently fragile, relying on code and requirements that may break between Linux, K8s, and containerd versions. In our opinion, the sheer amount of information required for gracefully terminating a download without disrupting any service, the inherent unpredictability of the system scheduler and the relative opacity of containerd’s download management render any further attempts in increasing MAGI’s speed better spent in improving the main problem instead.
@@ VII Related Work
Denial of Service based on Resources Usage. Resource-based DoS in Cloud Computing was previously covered in previous works. In [24], the authors present a novel class of attacks called Resource-Freeing Attacks (RFAs) and create a methodology to prevent the DoS by increasing the amount of resources available to the victim VM. Fang et al. also discussed resource and co-location attacks in [25] and [26], where they proposed respectively one new metric, called Heteroscore, and a threat method, REPTTACK. In the latter, a mitigation technique is also proposed, based on introducing randomness in the scheduling of the Virtual Machine (VM)s. However, both works do not consider container-based deployments, such as K8s. Finally, in [27], Zhan et al. show a framework that aims to detect CPU-exhaustion DoS attacks in containers called Coda. However, it does not consider the node resources, just the container ones. Finally, some CVEs related to resource-based DoS attacks in K8s have been reported, such as CVE-2022-1708161616CVE-2022-1708: https://nvd.nist.gov/vuln/detail/cve-2022-1708, and CVE-2025-0426171717CVE-2025-0426: https://nvd.nist.gov/vuln/detail/cve-2025-0426. Both CVEs deal with node resource exhaustion, the first one exploiting a CRI vulnerability causing large files to be written to disk, the second exploiting unauthenticated requests to cause a DoS condition in the node.
@@
Kubernetes and Container Security. Ahmed [28] shows that the container deployment process can generate significant resource usage, mainly memory and CPU. The author also tested several alterations to the container deployment’s decompress process, focusing on the parallelism of several simultaneous layers. Such approach decreases the total time to deploy a given image but inherently increases the resources used in this process.
Additionally, two studies investigate vulnerabilities in the configuration of various services within the Kubernetes orchestrator. He and Guo [3] show a threat in the resource configuration, where using highly privileged configurations allows eBPF applications to escape containers and access sensible information. Xiao [4], instead, found several threats based on the communication between microVMs solutions — such as Firecracker and Kata Containers — and containerd to escape attacks and resource-based DoS.
@@
Another important area that has received recent attention from the container orchestration community is anomaly detection in cluster resource usage. Almaraz-Rivera [5]’s study highlights the importance of implementing monitoring and alerting systems to track containers’ performance and resource utilization. He and Guo also implement a demo solution using an ML auto-encoder algorithm. Lastly, Yolchuyev [29] proposes an eXtreme Gradient Boosting (XGBoost) based model for anomaly detection on the Kubernetes orchestration platform. Either way, to the best of our knowledge, no prior work has been conducted on exploiting the asynchronous design of the K8s API communication.
Monitoring and Security in Kubernetes. While the academic community has reserved little attention in the field of monitoring and security in K8s, several tools in the market have been developed to address these issues.
OPA (Open Policy Agent) [30] is a tool that provides policy-based control for K8s. It allows users to define policies in a high-level language and enforce them before the K8s API server processes requests. For example, the use of the ---force flag can be actively blocked using OPA. However, OPA only performs its checks before the API server, which means that container image downloads cannot be controlled in this way.
Another popular tool for monitoring K8s environments is Falco [31] with its plugin k8saudit. In addition to its runtime monitoring capabilities, Falco can ingest K8s audit logs and generate alerts based on predefined rules. However, it works on a line-by-line basis and does not have built-in status tracking or counting features, which makes it unsuitable for tracking container image downloads. Using Falcosidekick [32] and a visualization tool like Grafana, this limitation can be overcome, but it requires additional setup and configuration efforts that may not be feasible in all environments. Still, Falco only performs monitoring and alerting, and does not provide any mitigation capabilities.
Finally, other tools such as Tetragon [33] and Tracee [34], provide runtime monitoring of containerized applications. Both tools are based on eBPF and can monitor system calls, network activity, and other events in real-time. They can also detect anomalous behavior and generate alerts based on predefined rules. Like Falco, both Tetragon and Tracee perform monitoring, and Tetragon is indeed actively used in MAGI to detect system calls. However, beyond their monitoring capabilities, they no means of actively mitigating the issues they detect.